OpenStudy (anonymous):

This is about volume of solid of revolutions. I have reviewed my practice problems and completed them again. Can someone please check two problems for me. They are posted below in comments.

OpenStudy (anonymous):

13. Volume = \[\int\limits_{0}^{6}[ \frac{ (-3)^{2}\pi \times6 }{ 3 }]dy\] 14. volume = \[\int\limits_{-4}^{6}[\pi(\frac{ -y-4 }{ 2 })^{2} - (3)^{2}]dx\]

OpenStudy (anonymous):

Please someone let me know if these are correct. I really want to make sure I understand this concept and I am doing these right.

OpenStudy (anonymous):

\[ \int_{-5}^{-2} 2 \pi (6-(-2 x-4)) (-x-2) \, dx=18 \pi \]

OpenStudy (anonymous):

This was for the first one

OpenStudy (accessdenied):

For (13), I observe that you used the volume of a cone formula with radius 3 and height 6 as the integrand. However, this already will give you the volume you were looking for without integration

OpenStudy (anonymous):

the problem, you put above, is that 14 and if so how did you get that? Would you know how to do thirteen with integration?

OpenStudy (anonymous):

For the second one \[ \int_{-5}^{-2} (6+3)^2 \pi \, dx-\int_{-5}^{-2} \pi ((-2 x-4)+3)^2 \, dx=243 \pi -117 \pi =126 \pi \]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8 this is the ones that I have not gotten. The other ones I had gotten help. These are the last two that im hoping I got right but through the comments above I don't think they are right. I don't know how those aove got to those answers though

ganeshie8 (ganeshie8):

can u go back to ur old post where we had drawings ?

OpenStudy (anonymous):

no problem! just comment on which one you want to use and ill go to it(:

ganeshie8 (ganeshie8):

ive tagged u there

ganeshie8 (ganeshie8):

il tag u again